# Scalable Multi-Objective Reinforcement Learning with Fairness Guarantees using Lorenz Dominance

### v1

@@ 8 Conclusion
We addressed key challenges in multi-objective, multi-policy reinforcement learning, by proposing methods that perform well over large state-action and objective spaces. We developed a new, multi-objective environment for simulating Public Transport Network Design, thereby enhancing the applicability of MORL to real-world scenarios. We proposed LCN, an adaptation of state-of-the-art methods that outperforms baselines in high-reward dimensions. Finally, we present an effective method for controlling the fairness constraint. These contributions move the research field toward more realistic and applicable solutions in real-world contexts, thereby advancing the state-of-the-art in algorithmic fairness in sequential decision-making and MORL.
Acknowledgments
@@
This research was in part supported by the European Union’s Horizon Europe research and innovation programme under grant agreement No 101120406 (PEER). DM is supported by the Innovation Center for AI (ICAI, The Netherlands). WR is supported by the Research Foundation – Flanders (FWO), grant number 1197622N. F. P. Santos acknowledges funding by the European Union (ERC, RE-LINK, 101116987).
A Theoretical Results
@@
In this section, we provide the detailed theoretical results for $\lambda$-LCN that we sketched on the main paper. In particular, we show that $\lambda$-Lorenz dominance is a generalisation of Lorenz dominance that can be used to flexibly set a desired fairness level. Moreover, by decreasing $\lambda$, the resulting solution set moves monotonically closer to the Lorenz front.
@@
${\bm{v}}\succ_{\lambda}{\bm{v}}^{\prime}\implies{\bm{v}}\succ_{L}{\bm{v}}^{\prime}.$ (13)
@@
${\bm{v}}\succ_{\lambda_{2}}{\bm{v}}^{\prime}\implies{\bm{v}}\succ_{\lambda_{1}}{\bm{v}}^{\prime}.$ (14)
@@
$\lambda_{2}\sigma({\bm{v}})_{i}+(1-\lambda_{2})L({\bm{v}})_{i}$ $\geq\lambda_{2}\sigma({\bm{v}}^{\prime})_{i}+(1-\lambda_{2})L({\bm{v}}^{\prime})_{i}$ (15)
$\frac{\lambda_{1}}{\lambda_{2}}\lambda_{2}\sigma({\bm{v}})_{i}+\frac{\lambda_{1}}{\lambda_{2}}(1-\lambda_{2})L({\bm{v}})_{i}$ $\geq\frac{\lambda_{1}}{\lambda_{2}}\lambda_{2}\sigma({\bm{v}}^{\prime})_{i}+\frac{\lambda_{1}}{\lambda_{2}}(1-\lambda_{2})L({\bm{v}}^{\prime})_{i}$ (16)
$\lambda_{1}\sigma({\bm{v}})_{i}+(\frac{\lambda_{1}}{\lambda_{2}}-\lambda_{1})L({\bm{v}})_{i}$ $\geq\lambda_{1}\sigma({\bm{v}}^{\prime})_{i}+(\frac{\lambda_{1}}{\lambda_{2}}-\lambda_{1})L({\bm{v}}^{\prime})_{i}$ (17)
$\lambda_{1}\sigma({\bm{v}})_{i}+(1-\lambda_{1})L({\bm{v}})_{i}$ $\geq\lambda_{1}\sigma({\bm{v}}^{\prime})_{i}+(1-\lambda_{1})L({\bm{v}}^{\prime})_{i}$ (18)
@@
${\bm{v}}\succ_{P}{\bm{v}}^{\prime}\implies{\bm{v}}\succ_{\lambda}{\bm{v}}^{\prime}.$ (19)
@@
From Definition 3, $\lambda$-Lorenz dominance can be written as $\lambda\sigma({\bm{v}})+(1-\lambda)L({\bm{v}})\succ_{P}\lambda\sigma({\bm{v}}^{\prime})+(1-\lambda)L({\bm{v}}^{\prime})$. Let us first recall that ${\bm{v}}\succ_{P}{\bm{v}}^{\prime}\implies{\bm{v}}\succ_{L}{\bm{v}}^{\prime}$ (?). It is then necessary to demonstrate an analogous result for ${\bm{v}}\succ_{P}{\bm{v}}^{\prime}\implies\sigma({\bm{v}})\succ_{P}\sigma({\bm{v}}^{\prime})$. We prove this by induction.
@@
The final case where $\sigma({\bm{v}})=({\bm{v}}_{2},{\bm{v}}_{1})$ and $\sigma({\bm{v}}^{\prime})=({\bm{v}}^{\prime}_{1},{\bm{v}}^{\prime}_{2})$, gives the same contradictions.
@@
Finally, we provide a proof for Theorem 1 of the main paper. This result demonstrates that when $\lambda$ is $1$, the solution set starts closest to the Pareto front and decreasing $\lambda$ to 00 monotonically reduces the solution set until it results in the Lorenz front. As such, by selecting a $\lambda$, a decision-maker can determine their preferred balance between Pareto optimality and Lorenz fairness. We first restate the theorem below and subsequently provide the proof.
@@
$\mathcal{L}\left(D\right)\subseteq\mathcal{L}\left(D;\lambda_{1}\right)\subseteq\mathcal{L}\left(D;\lambda_{2}\right)\subseteq\mathcal{F}\left(D\right).$ (20)
@@
B Preparing the Xi’an and Amsterdam Environments.
@@
We generated the Xi’an environment utilizing the data provided by ?. 333source: https://github.com/weiyu123112/City-Metro-Network-Expansion-with-RL. The city is divided into a grid of dimensions $H^{29\times 29}$, with cells of equal size ($1km^{2}$). The OD demand matrix was formulated using GPS data gathered from 25 million mobile phones, with their movements tracked over a one-month period Additionally, each cell is assigned an average house price index, which is categorized into quintiles. Figure 9 provides a comprehensive breakdown of the city into various sized groups.
@@
We generate and release the data associated with the Amsterdam environment. The city is divided into a grid of dimensions $H^{35\times 47}$, consisting of equally sizes cells of $0.5km^{2}$. The choice of this cell size takes into consideration Amsterdam’s smaller size compared to Xi’an. Since GPS data is unavailable for Amsterdam, we estimate the OD demand using the recently published universal law of human mobility, which states that the total mobility flow between two areas, denoted as $i$ and $j$, depends on their distance and visitation frequency (?). The estimation is computed using the formula:
$OD_{ij}=\mu_{j}\mathsf{K_{i}}/d^{2}_{ij}\ln(f_{max}/f_{min})$ (21)
@@
$\mu_{j}\approx\rho_{pop}(j)rad^{2}_{j}f_{max}$ (22)
Where $rad^{2}_{j}$ is the radius of area j. The flows are estimated for a full week, and in the model, this is accomplished by setting $f_{min}$ and $f_{max}$ to $1/7$ and $7$ respectively. Since the grid cells are of equal size in our case, the term $K$ can be omitted from the calculation. An illustration of the Amsterdam environment is presented in A detailed breakdown of the city into different-sized groups is depicted in Figure 10. Similar to the Xi’an environment, each cell in Amsterdam is associated with an average house price, sourced from the publicly available statistical bureau of The Netherlands dataset 444source: https://www.cbs.nl/nl-nl/maatwerk/2019/31/kerncijfers-wijken-en-buurten-2019.
Figure 9: MO-TNDP Xi’an Environment with different groups based on equally sized buckets of the average house price index.
Figure 10: MO-TNDP Amsterdam Environment with different groups based on equally sized buckets of the average house price index.
C Experiment Reproducibility Details
Each model presented in the paper for training MO-TNDP was trained for 30000 steps. Hyperparameter were tuned via a bayes search over 100 settings, with the following ranges:
@@ C Experiment Reproducibility Details
Figure 11: Architecture of the policy network.
Figure 12: Full results for the MO-TNDP Amsterdam and Xi’an Environments.
Figure 13: Learning curves for EUM.
Figure 14: Learning curves for Sen Welfare.
D
Additional Results
Gini coefficient: quantifies the reward distribution among various groups in the city. A value of 0 indicates perfect equality (equal percentage of satisfied OD flows per group), while 1 represents perfect inequality (only the demand of one group is satisfied). Although traditionally used to assess income inequality (?), it has also been employed in the context of transportation network design (?). In fig. 12 and table 2, we show detailed results on both Xi’an and Amsterdam for all objectives. fig. 13 and fig. 14 shows the learning curves of PCN and LCN for all objectives.
Table 2: Results of all models, for 1–10 objectives. Underline indicates the best results.
@@
GPI-LS $0.24\pm 0.08$ $0.13\pm 0.04$ $0.14\pm 0.04$ $0.01\pm 0.01$ $0.07\pm 0.04$-- - --- - --- - --- -
PCN $\underline{0.92\pm 0.02}$ $\underline{0.89\pm 0.03}$ $\underline{0.63\pm 0.10}$ $\underline{0.57\pm 0.11}$ $0.38\pm 0.09$ $0.11\pm 0.05$ $0.02\pm 0.01$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
LCN $0.81\pm 0.06$ $0.54\pm 0.02$ $0.62\pm 0.02$ $0.46\pm 0.07$ $0.50\pm 0.15$ $\underline{0.69\pm 0.08}$ $\underline{0.76\pm 0.07}$ $\underline{0.70\pm 0.09}$ $\underline{0.71\pm 0.16}$
LCN-Redist. $0.76\pm 0.03$ $0.44\pm 0.03$ $0.50\pm 0.06$ $\underline{0.57\pm 0.10}$ $\underline{0.86\pm 0.04}$ $0.44\pm 0.16$ $0.65\pm 0.05$ $0.39\pm 0.10$ $0.53\pm 0.22$
LCN-Mean $0.77\pm 0.03$ $0.43\pm 0.03$ $0.44\pm 0.03$ $0.32\pm 0.10$ $0.57\pm 0.04$ $0.32\pm 0.11$ $0.38\pm 0.06$ $0.36\pm 0.13$ $0.33\pm 0.13$
@@
GPI-LS $0.08\pm 0.04$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 --- - --- - --- - --- - --- -
PCN $\underline{0.92\pm 0.02}$ $\underline{0.95\pm 0.02}$ $0.21\pm 0.14$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
LCN $0.79\pm 0.05$ $0.65\pm 0.06$ $\underline{0.67\pm 0.17}$ $0.06\pm 0.06$ $\underline{0.20\pm 0.20}$ $\underline{0.23\pm 0.19}$ $\underline{0.39\pm 0.24}$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
LCN-Redist. $0.63\pm 0.08$ $0.53\pm 0.07$ $0.27\pm 0.17$ $\underline{0.31\pm 0.20}$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
LCN-Mean $0.75\pm 0.03$ $0.13\pm 0.08$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
@@
GPI-LS $0.32\pm 0.09$ $0.44\pm 0.12$ $0.46\pm 0.12$ $0.22\pm 0.08$ $0.57\pm 0.09$-- - --- - --- - --- -
PCN $\underline{0.90\pm 0.02}$ $\underline{0.97\pm 0.01}$ $\underline{0.90\pm 0.04}$ $\underline{0.95\pm 0.03}$ $\underline{0.84\pm 0.06}$ $0.66\pm 0.05$ $0.31\pm 0.11$ $\underline{0.52\pm 0.13}$ $0.19\pm 0.05$
LCN $0.85\pm 0.05$ $0.82\pm 0.01$ $0.74\pm 0.03$ $0.79\pm 0.02$ $0.64\pm 0.17$ $\underline{0.90\pm 0.02}$ $0.68\pm 0.10$ $0.49\pm 0.04$ $0.35\pm 0.07$
LCN-Redist. $0.83\pm 0.02$ $0.76\pm 0.02$ $0.63\pm 0.05$ $0.80\pm 0.04$ $\underline{0.87\pm 0.05}$ $0.77\pm 0.05$ $\underline{0.80\pm 0.12}$ $\underline{0.57\pm 0.05}$ $\underline{0.77\pm 0.09}$
LCN-Mean $0.85\pm 0.02$ $0.76\pm 0.02$ $0.62\pm 0.03$ $0.69\pm 0.05$ $0.70\pm 0.03$ $0.63\pm 0.17$ $0.20\pm 0.07$ $0.12\pm 0.06$ $0.21\pm 0.07$
@@
GPI-LS $0.21\pm 0.09$ $0.22\pm 0.11$ $0.28\pm 0.12$ $0.40\pm 0.13$-- - --- - --- - --- - --- -
PCN $\underline{0.94\pm 0.02}$ $\underline{0.94\pm 0.02}$ $0.78\pm 0.11$ $\underline{0.87\pm 0.04}$ $0.71\pm 0.07$ $0.55\pm 0.06$ $0.64\pm 0.07$ $0.27\pm 0.08$ $0.56\pm 0.09$
LCN $0.90\pm 0.03$ $0.78\pm 0.07$ $0.84\pm 0.03$ $0.66\pm 0.05$ $0.47\pm 0.05$ $0.46\pm 0.03$ $0.46\pm 0.06$ $0.42\pm 0.07$ $0.41\pm 0.04$
LCN-Redist. $0.79\pm 0.06$ $0.83\pm 0.06$ $\underline{0.89\pm 0.03}$ $\underline{0.89\pm 0.04}$ $\underline{0.95\pm 0.03}$ $\underline{0.90\pm 0.06}$ $\underline{0.82\pm 0.06}$ $\underline{0.69\pm 0.12}$ $\underline{0.96\pm 0.02}$
LCN-Mean $0.87\pm 0.02$ $0.50\pm 0.03$ $0.70\pm 0.06$ $0.56\pm 0.04$ $0.17\pm 0.06$ $0.15\pm 0.08$ $0.15\pm 0.06$ $0.15\pm 0.04$ $0.24\pm 0.07$
Normalized Total Efficiency
@@
GPI-LS $0.57\pm 0.02$ $0.48\pm 0.02$ $0.58\pm 0.01$ $0.23\pm 0.01$ $0.48\pm 0.01$-- - --- - --- - --- -
PCN $0.88\pm 0.01$ $0.90\pm 0.01$ $0.88\pm 0.00$ $\underline{0.82\pm 0.01}$ $0.82\pm 0.01$ $0.67\pm 0.01$ $0.61\pm 0.01$ $0.68\pm 0.01$ $0.63\pm 0.01$
LCN $0.87\pm 0.02$ $\underline{0.95\pm 0.01}$ $\underline{0.91\pm 0.01}$ $\underline{0.83\pm 0.03}$ $0.78\pm 0.04$ $\underline{0.83\pm 0.01}$ $\underline{0.92\pm 0.01}$ $0.91\pm 0.01$ $\underline{0.94\pm 0.01}$
LCN-Redist. $0.90\pm 0.01$ $0.92\pm 0.01$ $\underline{0.91\pm 0.01}$ $\underline{0.83\pm 0.01}$ $0.83\pm 0.01$ $0.76\pm 0.01$ $0.75\pm 0.01$ $0.69\pm 0.01$ $0.71\pm 0.01$
LCN-Mean $\underline{0.92\pm 0.01}$ $0.92\pm 0.02$ $0.86\pm 0.01$ $0.87\pm 0.02$ $\underline{0.96\pm 0.01}$ $0.79\pm 0.01$ $0.84\pm 0.01$ $\underline{0.93\pm 0.01}$ $0.86\pm 0.02$

### v2

@@ 8. Conclusion
We addressed key challenges in multi-objective, multi-policy reinforcement learning by proposing methods that perform well over large state-action and objective spaces. We developed a new multi-objective environment for simulating Public Transport Network Design, thereby enhancing the applicability of MORL to real-world scenarios. We proposed LCN, an adaptation of state-of-the-art methods that outperforms baselines in high-reward dimensions. Finally, we present an effective method for controlling the fairness constraint. These contributions move the research field toward more realistic and applicable solutions in real-world contexts, thereby advancing the state-of-the-art in algorithmic fairness in sequential decision-making and MORL.
Acknowledgements.
@@
This research was in part supported by the European Union’s Horizon Europe research and innovation program under grant agreement No 101120406 (PEER). DM is supported by the Innovation Center for AI (ICAI, The Netherlands). WR is supported by the Research Foundation – Flanders (FWO), grant number 1197622N. F. P. Santos acknowledges funding by the European Union (ERC, RE-LINK, 101116987).
Appendix A Theoretical Results
@@
In this section, we provide the detailed theoretical results for $\lambda$-LCN that we sketched in the main paper. In particular, we show that $\lambda$-Lorenz dominance is a generalization of Lorenz dominance that can be used to flexibly set a desired fairness level. Moreover, by decreasing $\lambda$, the resulting solution set moves monotonically closer to the Lorenz front.
@@
(13) ${\bm{v}}\succ_{\lambda}{\bm{v}}^{\prime}\implies{\bm{v}}\succ_{L}{\bm{v}}^{\prime}.$
@@
(14) ${\bm{v}}\succ_{\lambda_{2}}{\bm{v}}^{\prime}\implies{\bm{v}}\succ_{\lambda_{1}}{\bm{v}}^{\prime}.$
@@
(15) $\lambda_{2}\sigma({\bm{v}})_{i}+(1-\lambda_{2})L({\bm{v}})_{i}$ $\geq\lambda_{2}\sigma({\bm{v}}^{\prime})_{i}+(1-\lambda_{2})L({\bm{v}}^{\prime})_{i}$
(16) $\frac{\lambda_{1}}{\lambda_{2}}\lambda_{2}\sigma({\bm{v}})_{i}+\frac{\lambda_{1}}{\lambda_{2}}(1-\lambda_{2})L({\bm{v}})_{i}$ $\geq\frac{\lambda_{1}}{\lambda_{2}}\lambda_{2}\sigma({\bm{v}}^{\prime})_{i}+\frac{\lambda_{1}}{\lambda_{2}}(1-\lambda_{2})L({\bm{v}}^{\prime})_{i}$
(17) $\lambda_{1}\sigma({\bm{v}})_{i}+(\frac{\lambda_{1}}{\lambda_{2}}-\lambda_{1})L({\bm{v}})_{i}$ $\geq\lambda_{1}\sigma({\bm{v}}^{\prime})_{i}+(\frac{\lambda_{1}}{\lambda_{2}}-\lambda_{1})L({\bm{v}}^{\prime})_{i}$
(18) $\lambda_{1}\sigma({\bm{v}})_{i}+(1-\lambda_{1})L({\bm{v}})_{i}$ $\geq\lambda_{1}\sigma({\bm{v}}^{\prime})_{i}+(1-\lambda_{1})L({\bm{v}}^{\prime})_{i}$
@@
(19) ${\bm{v}}\succ_{P}{\bm{v}}^{\prime}\implies{\bm{v}}\succ_{\lambda}{\bm{v}}^{\prime}.$
@@
From Definition 3, $\lambda$-Lorenz dominance can be written as $\lambda\sigma({\bm{v}})+(1-\lambda)L({\bm{v}})\succ_{P}\lambda\sigma({\bm{v}}^{\prime})+(1-\lambda)L({\bm{v}}^{\prime})$. Let us first recall from Theorem 1 of (perny_approximation_2013) that ${\bm{v}}\succ_{P}{\bm{v}}^{\prime}\implies{\bm{v}}\succ_{L}{\bm{v}}^{\prime}$. It is then necessary to demonstrate an analogous result for ${\bm{v}}\succ_{P}{\bm{v}}^{\prime}\implies\sigma({\bm{v}})\succ_{P}\sigma({\bm{v}}^{\prime})$. We prove this by induction.
@@
The final case, where $\sigma({\bm{v}})=({\bm{v}}_{2},{\bm{v}}_{1})$ and $\sigma({\bm{v}}^{\prime})=({\bm{v}}^{\prime}_{1},{\bm{v}}^{\prime}_{2})$, gives the same contradictions.
@@
Finally, we provide a proof for Theorem 1 of the main paper. This result demonstrates that when $\lambda$ is $1$, the solution set starts closest to the Pareto front and decreasing $\lambda$ to 0 monotonically reduces the solution set until it results in the Lorenz front. As such, by selecting a $\lambda$, a decision-maker can determine their preferred balance between Pareto optimality and Lorenz fairness. We first restate the theorem below and subsequently provide the proof.
@@
(20) $\mathcal{L}\left(D\right)\subseteq\mathcal{L}\left(D;\lambda_{1}\right)\subseteq\mathcal{L}\left(D;\lambda_{2}\right)\subseteq\mathcal{F}\left(D\right).$
@@
Appendix B Preparing the Xi’an and Amsterdam Environments
@@
We generated the Xi’an environment utilizing the data provided in (wei_city_2020). 333source: https://github.com/weiyu123112/City-Metro-Network-Expansion-with-RL. The city is divided into a grid of dimensions $H^{29\times 29}$, with cells of equal size ($1km^{2}$). The OD demand matrix was formulated using GPS data gathered from 25 million mobile phones, with their movements tracked over a one-month period Additionally, each cell is assigned an average house price index, which is categorized into quintiles. Figure 10 provides a comprehensive breakdown of the city into various sized groups.
@@
We generate and release the data associated with the Amsterdam environment. The city is divided into a grid of dimensions $H^{35\times 47}$, consisting of equally sizes cells of $0.5km^{2}$. The choice of this cell size takes into consideration Amsterdam’s smaller size compared to Xi’an. Since GPS data is unavailable for Amsterdam, we estimate the OD demand using the recently published universal law of human mobility, which states that the total mobility flow between two areas, denoted as $i$ and $j$, depends on their distance and visitation frequency (schlapfer_universal_2021). The estimation is computed using the formula:
(21) $OD_{ij}=\mu_{j}\mathsf{K_{i}}/d^{2}_{ij}\ln(f_{max}/f_{min})$
@@
(22) $\mu_{j}\approx\rho_{pop}(j)rad^{2}_{j}f_{max}$
Where $rad^{2}_{j}$ is the radius of area j. The flows are estimated for a full week, and in the model, this is accomplished by setting $f_{min}$ and $f_{max}$ to $1/7$ and $7$ respectively. Since the grid cells are of equal size in our case, the term $K$ can be omitted from the calculation. An illustration of the Amsterdam environment is presented in A detailed breakdown of the city into different-sized groups is depicted in Figure 11.
Similar to the Xi’an environment, each cell in Amsterdam is associated with an average house price, sourced from the publicly available statistical bureau of the Netherlands dataset 444source: https://www.cbs.nl/nl-nl/maatwerk/2019/31/kerncijfers-wijken-en-buurten-2019.
Map of the MO-TNDP Xi’an environment showing different groups based on equally sized buckets of the average house price index. Each group represents a range of house prices.
Figure 10. MO-TNDP Xi’an Environment with different groups based on equally sized buckets of the average house price index.
Map of the MO-TNDP Amsterdam environment showing different groups based on equally sized buckets of the average house price index. Each group represents a range of house prices.
Figure 11. MO-TNDP Amsterdam Environment with different groups based on equally sized buckets of the average house price index.
Appendix C Experiment Reproducibility Details
Each model presented in the paper for training MO-TNDP was trained for 30000 steps. Hyperparameters were tuned via a Bayes search over 100 settings, with the following ranges:
@@ Appendix C Experiment Reproducibility Details
Diagram showing the architecture of the policy network, including input features, the state and condition embedding, two linear layers after the embedding, and the action probability output.
Figure 12. Architecture of the policy network.
Figure 13. Sensitivity analysis of the crowding-distance penalty: Xi’an Environment, 2 groups (A, B, C) and 5 groups (D, E, F). (A) Randomly sampled experiences, with the non-dominated set and the LCN-Mean and LCN-Redist reference points highlighted. (B) Fraction of sampled experiences that receive a crowding penalty as a function of the crowding-distance threshold ($\tau_{cd}$).(C) Kernel density estimates (KDEs) of the total penalty applied to the experiences for different penalty multipliers, assuming $\tau_{cd}=0.2$.
Appendix D Sensitivity Analysis of Crowding Distance Threshold and Penalty
We conduct a sensitivity analysis on the effects of the crowding-distance threshold $\tau_{cd}$ and the penalty multiplier $\rho_{pen}$. When the model filters experiences to maintain a consistent Experience Replay (ER) buffer, it applies a distance metric together with a crowdedness threshold. Experiences that lie too close to each other, i.e., those whose crowding distance exceeds the threshold, are penalized to promote diversity, as the objective is to train a broad set of policies for the decision-maker. We examine, in a theoretical setting, how variations in $\tau_{cd}$ and $\rho_{pen}$ influence both the proportion of penalized samples and the overall performance of the algorithm.
In Figure 13, we generated random sets of experiences in a 2-dimension (Panel A) and 5-dimension setting (Panel D). We show the non-dominated points, as well as the LCN-Mean and LCN-Redist reference points. In Panels C and E we show the sensitivity of the generated experiences to the crowding distance threshold.
The influence of the crowding distance threshold ($\tau_{cd}$) varies significantly with the dimensionality of the objective space. In low dimensions (e.g., 2D), experiences cluster closer in the crowding-distance metric. Consequently, the sensitivity to $\tau_{cd}$ is high initially, with the proportion of penalized points peaking sharply at a very low threshold (e.g., $\tau_{cd}\approx 0.1$), where nearly all points are penalized. This calls for setting a low $\tau_{cd}$ in low-dimensional settings for the crowding distance penalty to remain effective. Conversely, in higher dimensions (e.g., 5D), the crowding distance penalty peaks at a later stage (e.g., $\tau_{cd}\approx 0.2$). The points are naturally more spread out, giving the designer finer control over the number of penalized experiences.
The penalty multiplier ($\rho_{pen}$) impacts the distribution of the penalty-adjusted distance scores. With a low penalty multiplier ($\rho_{pen}=1$), the distribution of the adjusted scores is concentrated close to zero, with low variance and thin tails, leading to a minimal scaling effect beyond the crowding distance. However, increasing $\rho_{pen}$ significantly skews the distribution, introducing heavier tails and increasing the variance of the penalty. This high multiplier aggressively penalizes points below the threshold, meaning points just below the threshold are greatly favored over those further away. A higher $\rho_{pen}$ is desirable in low-diversity environments (where policies are very close) to discriminate between them. A lower $\rho_{pen}$ suffices in high-diversity environments where distance itself is a sufficient discriminator. In this paper, we scale our environment to high dimensions, however, we operate in a relatively low-variance environment with hard constraints on the generated policies, hence we opted for a middle-point penalty of $\rho_{pen}=2$.
Appendix E Additional Results
Comprehensive results for the MO-TNDP Amsterdam and Xi’an environments.
Figure 14. Full results for the MO-TNDP Amsterdam and Xi’an Environments.
Learning curves for the EUM metric across multiple objectives, showing convergence over training iterations.
Figure 15. Learning curves for EUM.
Learning curves for the Sen Welfare metric across multiple objectives.
Figure 16. Learning curves for Sen Welfare.
Table 2. Results of all models, for 1–10 objectives. Underline indicates the best results.
@@ Appendix E Additional Results
GPI-LS $11\pm 0.03 0.0$$8\pm 0.02 0.0$$5\pm 0.03 0.0$$3\pm 0.01 0.0$$02\pm 0.01 -- -- -- --$
PCN $\underline{0.89\pm 0.02}$ $\underline{0.93\pm 0.02}$ $\underline{0.83\pm 0.04}$ $0.63\pm 0.08$ $0.56\pm 0.08$ $17\pm 0.04 0.0$$7\pm 0.03 0.0$$2\pm 0.02 0.0$$00\pm 0.00$
LCN $0.86\pm 0.03$ $0.71\pm 0.03$ $0.54\pm 0.04$ $\underline{0.71\pm 0.06}$ $\underline{0.67\pm 0.06}$ $\underline{0.55\pm 0.08}$ $\underline{0.60\pm 0.08}$ $0.29\pm 0.11$ $\underline{0.46\pm 0.07}$
LCN-Redist. $0.79\pm 0.01$ $0.69\pm 0.03$ $0.48\pm 0.06$ $0.60\pm 0.04$ $0.64\pm 0.05$ $0.39\pm 0.03$ $0.53\pm 0.08$ $\underline{0.31\pm 0.11}$ $0.19\pm 0.08$
LCN-Mean $0.78\pm 0.02$ $0.70\pm 0.02$ $0.52\pm 0.07$ $0.51\pm 0.05$ $0.49\pm 0.08$ $0.37\pm 0.08$ $0.12\pm 0.08$ $0.15\pm 0.08$ $0.28\pm 0.10$
@@
GPI-LS $0.12\pm 0.05$ $22\pm 0.06 0.0$$6\pm 0.06 0.0$$00\pm 0.00 -- -- -- -- --$
PCN $\underline{0.97\pm 0.01}$ $\underline{0.80\pm 0.04}$ $17\pm 0.11 0.0$$\pm 0.00 0.00±0$$\pm 0.00 0.0$$\pm 0.00 0.00±0$$\pm 0.00 0.0$$\pm 0.00 0.00±0$$0\pm 0.00$
LCN $0.80\pm 0.03$ $0.67\pm 0.04$ $\underline{0.41\pm 0.14}$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $\pm 0.00 0.00±0.00$$\pm 0.00$
LCN-Redist. $0.79\pm 0.03$ $53\pm 0.03 0.0$$\pm 0.00 0.00±0$$\pm 0.00 0.0$$\pm 0.00 0.00±0$$\pm 0.00 0.0$$\pm 0.00 0.00±0$$\pm 0.00 0.0$$00\pm 0.00$
LCN-Mean $0.81\pm 0.01$ $27\pm 0.05 0.0$$\pm 0.00 0.00±0$$\pm 0.00 0.0$$\pm 0.00 0.00±0$$\pm 0.00 0.0$$\pm 0.00 0.00±0$$\pm 0.00 0.0$$00\pm 0.00$
@@
GPI-LS $0.27\pm 0.06$ $0.24\pm 0.06$ $0.33\pm 0.08$ $0.42\pm 0.08$ $0.38\pm 0.08$ -- -- -- --
PCN $\underline{0.96\pm 0.01}$ $0\underline{.97\pm 0.01}$ $\underline{0.94\pm 0.02}$ $0.86\pm 0.10$ $0.83\pm 0.04$ $\underline{0.57\pm 0.09}$ $0.55\pm 0.04$ $0.39\pm 0.07$ $0.27\pm 0.05$
LCN $0.95\pm 0.01$ $0.89\pm 0.02$ $0.79\pm 0.02$ $\underline{0.91\pm 0.02}$ $0.80\pm 0.02$ $0.50\pm 0.06$ $\underline{0.74\pm 0.03}$ $0.42\pm 0.04$ $0.38\pm 0.04$
LCN-Redist. $0.92\pm 0.01$ $0.88\pm 0.02$ $0.73\pm 0.04$ $0.89\pm 0.02$ $\underline{0.83\pm 0.04}$ $0.52\pm 0.07$ $0.63\pm 0.07$ $\underline{0.86\pm 0.05}$ $\underline{0.92\pm 0.02}$
LCN-Mean $0.91\pm 0.01$ $0.89\pm 0.01$ $0.78\pm 0.04$ $0.84\pm 0.02$ $0.71\pm 0.03$ $0.26\pm 0.09$ $0.39\pm 0.07$ $5\pm 0.05 0.22±0$$2\pm 0.06$
@@
GPI-LS $0.30\pm 0.07$ $0.85\pm 0.05$ $0.92\pm 0.02$ $0.67\pm 0.07$ -- -- -- -- --
PCN $\underline{0.99\pm 0.00}$ $\underline{0.95\pm 0.01}$ $\underline{0.75\pm 0.04}$ $\underline{0.80\pm 0.04}$ $0.76\pm 0.04$ $0.62\pm 0.09$ $0.59\pm 0.07$ $0.36\pm 0.10$ $0.32\pm 0.09$
LCN $0.93\pm 0.02$ $0.70\pm 0.03$ $0.21\pm 0.05$ $0.55\pm 0.02$ $0.58\pm 0.02$ $0.47\pm 0.04$ $0.79\pm 0.02$ $0.44\pm 0.03$ $0.50\pm 0.03$
LCN-Redist. $0.92\pm 0.02$ $0.55\pm 0.08$ $0.46\pm 0.07$ $0.79\pm 0.02$ $\underline{0.86\pm 0.02}$ $\underline{0.89\pm 0.03}$ $\underline{0.90\pm 0.02}$ $\underline{0.88\pm 0.02}$ $\underline{0.92\pm 0.01}$
LCN-Mean $0.94\pm 0.01$ $0.15\pm 0.06$ $0.21\pm 0.05$ $0.17\pm 0.08$ $2\pm 0.07 0.11±0$$1\pm 0.02$ $0.12\pm 0.03$ $0.22\pm 0.01$ $0.36\pm 0.02$
Normalized Sen Welfare
@@
GPI-LS $0.21\pm 0.04$ $0.26\pm 0.02$ $0.23\pm 0.01$ $0.25\pm 0.01$ $0.20\pm 0.01$ -- -- -- --
PCN $0.87\pm 0.01$ $0.78\pm 0.01$ $0.67\pm 0.01$ $0.71\pm 0.01$ $0.64\pm 0.01$ $0.51\pm 0.01$ $0.46\pm 0.01$ $0.45\pm 0.01$ $0.43\pm 0.00$
LCN $0.93\pm 0.01$ $0.91\pm 0.02$ $0.86\pm 0.02$ $\underline{0.91\pm 0.00}$ $\underline{0.87\pm 0.01}$ $\underline{0.79\pm 0.01}$ $\underline{0.77\pm 0.01}$ $0.62\pm 0.02$ $\underline{0.81\pm 0.01}$
LCN-Redist. $\underline{0.97\pm 0.01}$ $\underline{0.96\pm 0.01}$ $0.82\pm 0.02$ $7\pm 0.01 0.77±0$$7\pm 0.01$ $0.62\pm 0.01$ $0.59\pm 0.01$ $0.54\pm 0.01$ $\underline{0.56\pm 0.01}$
LCN-Mean $0.90\pm 0.01$ $0.92\pm 0.02$ $\underline{0.87\pm 0.01}$ $0.90\pm 0.01$ $0.84\pm 0.01$ $0.78\pm 0.01$ $0.71\pm 0.01$ $\underline{0.76\pm 0.01}$ $0.80\pm 0.01$